OpenStudy (toxicsugar22):

can anyone help me with this one (2x-1)^2=3

OpenStudy (anonymous):

2x-1 = -sqrt(3) and 2x-1 =sqrt(3) solve for x

OpenStudy (toxicsugar22):

can u help me do that

OpenStudy (anonymous):

Help to solve for x?

OpenStudy (toxicsugar22):

yes

OpenStudy (toxicsugar22):

solve the equation

OpenStudy (anonymous):

2x -1 = sqrt(3) 2x = sqrt(3) + 1 x = (sqrt(3)+1)/2 You can apply the same approach for the second part

OpenStudy (toxicsugar22):

so is the anwer 1+ or - sqrt3/2

OpenStudy (anonymous):

x= \[ \frac{(\sqrt{3} +1)}{2} \]

OpenStudy (toxicsugar22):

OK THANKS

OpenStudy (anonymous):

You need to solve the second solution for 2x-1 = -sqrt(3) using the steps I show above. http://www.oojih.com for free study guide and free practice test

OpenStudy (toxicsugar22):

thanks